. An electron of mass me, initially at rest, moves though a certain distance in a uniform electric field in time t2 A proton of mass mp, aslo, initially at rest, takes time t to move though an equal distance in this uniform electric field. Neglecting the effect of gravity, the ratio t2 / t1 is nearly equal to
 
(a) 1
(b) (mp / me)1/2
(c) (me / mp)1/2
(d) 1836

Grade:11

1 Answers

Kevin Nash
9 years ago
Hello Student,
Please find the answer to your question
. (b) Let the distance to be travelled be x, Let the strength of uniform electric field be E, For electron
u = 0, s = x, a = eE/me , t = t1 , S = ut + ½ eE/2e x t21
⇒ x = ½ eE/2 me x t21 …..(i)
For proton
u = 0, s = x, a = eE/ mp , t = t2, S = ut + ½ at2
⇒ x = ½ eE/2mp x t22 ……(ii)
From (i) and (ii)
t22 / t21 = mp/me ⇒ t2/ t1 = [mp / me]1/2
Thanks
Kevin Nash
askIITians Faculty